OpenStudy (anonymous):

(y+7)^2=-3 solve by the square root method...please help

OpenStudy (anonymous):

im assuming you're solving for y?

OpenStudy (anonymous):

(y+7)^2 = (y+7)(y+7) <-- FOIL y^2+14y+49= -3 y^2 + 14y + 52 = 0 and i just got stuck gimme a sec lol try to refresh my memory

OpenStudy (anonymous):

and just to clarify are you sure you're supposed to use the square root method on that problem?

OpenStudy (anonymous):

yeah that's what the instructions said to do

OpenStudy (anonymous):

i just looked it up and it says to sqrt both sides, but if you sqrt a negative number you're left with i. have you learned those yet?

OpenStudy (anonymous):

You should take sqrt in both sides but -3 that will be unreal number. Sqrt(-3)=+,- i sqrt(3) Therefore you will get: y+7=+,- i sqrt(3) y=-7+,- i sqrt(3)

OpenStudy (anonymous):

\[(y+7)(y+7)=-3\] \[y^2+14y+49=-3\] \[y^2+14y+52=0\] use quadratic from here it easier \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] \[\frac{ -14\pm \sqrt{14^2-4(1)(52)} }{ 2(1) }\] \[\frac{ -14\pm \sqrt{196-208} }{ 2 }\] \[\frac{ -14\sqrt{-12} }{ 2 }\] \[\frac{ -14\pm2i \sqrt{3} }{ 2 }\] \[\frac{ -14+2i \sqrt{3} }{ 2 } and \frac{ -14-2i \sqrt{3} }{ 2}\]

OpenStudy (anonymous):

pretty sure she is required to use the square root method

OpenStudy (anonymous):

and it seems like you get the same answer anyway with a lot less work, you can simplify both sides by 2 to get -7 - isqrt 3

OpenStudy (anonymous):

actually make sure its +/- thats a big part of it

OpenStudy (anonymous):

ok thank guys